OpenStudy (alante):

Somebody Help Johnny has 15 coins in his pocket, consisting of nickels and dimes. The total value of the coins is $1.15. How many of each coin does Johnny have? (Write the nickels first then the dimes, just type the numbers i.e. 15, 7 )

OpenStudy (sagewilson):

Okay I got it, it took me some time. You would need to write 7, 8. If you had 8 dimes and 7 nickles: 5 * 7 = 35 (0.35) 10 * 8 = 80 (.80) If you added those together you would get 1.15.

OpenStudy (alante):

Thats It? @SageWilson

OpenStudy (sagewilson):

Yup, (7, 8), that's your answer.

OpenStudy (anonymous):

Well that is definitely not the method you should use, but it is the answer.

OpenStudy (sagewilson):

In all honesty I wasn't exactly sure what method I should have used so I just ended up typing in numbers into a calculator until I came up with 1.15.

OpenStudy (anonymous):

This is fairly obviously a system of equations problem.